General Meeting Summary 7/21/22

From Cuawiki
Jump to: navigation, search

PARTICIPANTS: Hamlet Mkrtchyan, Yeranuhi Gavilyan, Brad Sawatzky, Arshak Asaturyan, Vladimir Berdnikov, Tanja Horn, Carlos Munoz-Camacho, Fernando Barbosa, Paulo Medeiros, Ye, Zhenyu, Rolf Ent, Abishek Karki, Charles Hyde, Pramita Tiwari, Vardan Tadevosyan


NPS ASSEMBLY - INITIAL PMT CHECKOUT (Brad, Pramita, Fernando, Chris, et al.)

  • put together baseline software and hardware systems and able to take a first look at a couple NPS PMTs
  • needed to reconfigure the MPOD module for the slow regulation for the 150 foot cable.
  • There seems to be a problem with at least 1 of the 4 connected LV distribution boards.
  • Focusing on the first column of cells, the current draw for that LV board matched the 18mA/ base prediction and everything behaved reasonably well.
  • Ramped up HV on one PMT while watching the scope with a small effective threshold (relative to the noted baseline). Current

draw on the HV supply at 600V was 400--405 uA -- exactly what was projected.

  • Saw ~30mV level cosmics pulses on the 3 PMTs looked at with the HV at 700V. This is consistent with what wase expected from Vlad's studies.
  • There is some significant 20--30mV level, high frequency noise present on the baseline right now. Grounding is not yet optimal, so:
  • will go through the other LV boards/connections tomorrow and locate what is driving the higher than expected current draw.
  • operate PMTs at up to 800 V to pull out the noise - should be safe for the 8-stage biasing for tests. For normal operation, HV=605 @ 405 uA, nominally, to limit the anode current
  • Action items:
  • understand noise and try to reduce/eliminate it
  • plan is to finish the setup for the four columns by end of the month


ELECTRONICS

  • Major item being addressed is event unblocking
  • Diagnostic tools will be available for detector checkout - based on existing tools and adapted by Pramita/Julie


DSG UPDATES (Aaron)

  • High Voltage Settings GUI - allows for setting parameters for all CAEN modules/channels
  • EPICS screens development for monitoring - debugging and optimizations
  • Chiller testing - communication and flow meter testing; plan to start next week
  • Started process of installing applications for monitoring etc. on Hall C computers, but encountered difficulties


NPS MECHANICAL (Paulo)

  • Detector lift fixtures and rail section
  • presently being painted and will be delivered by the end of the month
  • Cable carrier/guide evaluation
  • cable bundles on dollies seems to be the best option for NPS - carrier might become to large and heavy
  • NPS beam pipe 1 out for procurement
  • beam-line 2 is under physics review


NPS PLANNING

  • Plan for cabling and PMT final assembly etc.
  • IJCLab-Orsay team plans to arrive in second half of September and team members will be on site for 1-1.5 months to complete the tasks and possibly contribute to detector checkout
  • helpful to have initial PMT assembly ready
  • Initial PMT assembly (AANL team, Vladimir)
  • AANL team will lead the initial PMT assembly as soon as the space is available - need a table - perhaps Mezzanine would be a good place?
  • additional contributors welcome! Please contact Tanja/Brad/Vladimir
  • Near term cabling needs to be complete
  • presently the team responsible is: AANL team, Pramita, Erik, Abishek
  • Action item:
  • update Gantt chart looking forward to the next year
  • include bullet points with tasks (also include administrative steps towards the Phase 1 experiments like any needed review)
  • PAC50
  • DVCS on neutron approved - next steps: contact Patrizia if ERR needed and submit beam time request
  • TCS experiment deferred for opportunity to demonstrate experiment design

NEXT MEETING: THURSDAY 25 AUGUST AT 9AM ET